Pamela is 9 years older than Jiri. The sum of their ages is 107 . What is Jiri's age?
Answer by london maths tutor(243)

Let's Jiri's age be "J" and Pamela's age be "P".
P = 9 + J (Pamela is 9 years older than Jiri) ... equation 1
P + J = 107 (the sum of their ages is 107) ... equation 2
Put equation 1 into equation 2:
(9+J) + J = 109
9 + 2J = 109
2J = 109 - 9
2J = 100
J = 100/2
J = 50
Therefore,Jiri's age is 50.